OpenStudy (anonymous):

The first term of an arithmetic sequence is -3 and the fifteenth term is 53. What is the common difference of the sequence?

OpenStudy (anonymous):

53 - (-3) = 56 that's the total difference from the first term to the 15th term. So the common difference is 56/(15-1) (starting pos - ending pos) = 56/14 = 4

OpenStudy (anonymous):

thank you!